# Benchmarking Unsupervised Online IDS for Masquerade Attacks in CAN

### v2

@@ Abstract
Vehicular controller area networks (CANs) are susceptible to masquerade attacks by malicious adversaries. In masquerade attacks, adversaries silence a targeted ID and then send malicious frames with forged content at the expected timing of benign frames. As masquerade attacks could seriously harm vehicle functionality and are the stealthiest attacks to detect in CAN, recent work has devoted attention to compare frameworks for detecting masquerade attacks in CAN. However, most existing works report offline evaluations using CAN logs already collected using simulations that do not comply with the domain’s real-time constraints. Here we contribute to advance the state of the art by introducing a benchmark study of four different non-deep learning (DL)-based unsupervised online intrusion detection systems (IDS) for masquerade attacks in CAN. Our approach differs from existing benchmarks in that we analyze the effect of controlling streaming data conditions in a sliding window setting. In doing so, we use realistic masquerade attacks being replayed from the ROAD dataset. We show that although benchmarked IDS are not effective at detecting every attack type, the method that relies on detecting changes in the hierarchical structure of clusters of time series produces the best results at the expense of higher computational overhead. We discuss limitations, open challenges, and how the benchmarked methods can be used for practical unsupervised online CAN IDS for masquerade attacks.
Index Terms:
Streaming Data, Online Algorithms, Machine Learning, Anomaly Detection, In-Vehicle Security, CAN Bus.
I Introduction
@@
IDS for detecting masquerade attacks in CAN are often based on unsupervised anomaly detection techniques aiming to characterize the regular relationships between physical signals in a vehicle [15, 18]. Some of them focus on mining time series relationships based on deep learning (DL) models that consequently have high computational cost [19, 16, 20]. In addition, most existing CAN IDS approaches are tested offline using datasets or CAN logs already collected from a real or simulated environment. This has been synthesized in existing benchmarking frameworks and comparative studies in the CAN IDS space [21, 18]. The offline evaluation approach is different from online evaluation, in that the latter is performed using streaming CAN data from a vehicle, simulation, or data log replay [22]. Therefore, offline evaluation hinders the understanding of realistic capabilities and limitations of CAN IDS in a streaming environment. By benchmarking non DL-based unsupervised online IDS for masquerade attacks in CAN, we seek to provide insights into the effectiveness and limitations of these methods in resource-constrained and streaming environments.
@@
In this paper, we address this gap of knowledge by introducing an empirical and unbiased benchmark of four different non-DL-based unsupervised online IDS for masquerade attacks: Matrix Collelation Distribution, Matrix Correlation Correlation, Ganesan17 [23], and Moriano22 [15]. We focus on non DL-based approaches because vehicular data change over time and space. This dynamic nature makes it difficult for DL-based methods to extract relevant features, often leading to poor performance [24]. Here, Matrix Correlation Distribution and Matrix Correlation Correlation methods are simple proposed baselines while Ganesan17 and Moriano22 are already proposed methods in the literature. We test them in simulated masquerade attacks from the ROAD dataset [10]. We rely on a sliding window approach to simulate data log replays of a continuous data stream [25, 26]. We analyze the joint effect of window size and offset on the AUC-ROC as evaluation metric for intrusion detection across the different attacks in the ROAD dataset. We also report an analysis of the testing time per window (TTW) [20] of CAN IDS as a proxy of inference time and discuss insights into the real-time detection capability of these approaches. We further contribute to advance the state of the art in this topic by making publicly available the code of the IDS considered in this study [27]. This furthers research in this topic by allowing other researchers to replicate our results and compare new algorithms with the benchmarked IDS.
@@
Our results indicate that overall the benchmarked unsupervised online CAN IDS are not effective in consistently detecting every attack type in the ROAD dataset. Attacks in the ROAD dataset have different levels of impact on the regular correlation patterns of signals. However, evaluation performance metrics tend to favor Moriano22 and Matrix Correlation Distribution methods. This aligns with the expectation that impacting correlations across signals tends to disrupt the hierarchy of their relationships and their pairwise similarity distribution. Notably, among the variety of attacks in the ROAD dataset, max_engine_attack and correlated_attack are the ones that can be detected more easily across CAN IDS. Our results also shed light on the impact that sliding window parameters (i.e., window length and offset) have on the detection task. Overall, across methods and attack types, we identify conditions that make certain algorithms more suitable to operate in a real-time detection environment.
II Background
II-A CAN Protocol
@@ II-C CAN Vulnerabilities
  • Masquerade attacks: This advanced attack requires two compromised ECUs, one as a strong attacker and one as a weak attacker. The weak attacker monitors the frequency of a targeted ID and stops the legitimate ECU once it learns the frequency. The strong attacker then injects messages with the same ID at the original frequency. The ID frequency remains unchanged, but the source changes, and the messages contain forged content.
@@
III Related Work
Here we review and contrast previous offline benchmark studies on CAN IDS (Section III-A) and online CAN IDS methods (Section III-B).
III-A Prior Work Closely Related to the Present Study
Ji et al. [21] presented a comparative analysis of four statistical-based CAN IDSes using entropy, clock skew, ID sequences, and CAN bus throughput. Benchmarked methods relied only on ID timestamp arrivals. Their evaluation was done in simulated attack datasets. They found that the clock skew method performed best across attacks while the entropy method falls short on detecting replay attacks.
Dupont et al. [36] introduced a unifying framework for CAN IDS evaluation of time-based methods. In doing so, they used data from two live vehicles and a CAN prototype. They used both simulated attacks and publicly available datasets with attacks [37]. They found that among compared methods, they perform well only on attacks that generate significant disruptions in CAN traffic. They suggested that methods that consume bit representations and semantics of CAN messages will perform better.
Stachowski et al. [38] introduced a framework for benchmarking CAN IDS. Their approach focused on the online evaluation of three existing solutions from identified vendors, so that CAN IDS are tested in real vehicles for real-time intrusion detection. Targeted ID attacks were introduced in test vehicles while vehicles were in motion and stationary. None of the tested CAN IDS was found effective for detecting the variety of attacks.
Blevins et al. [35] focused on bechmarking time-based CAN IDS on the fuzzy and targetted ID attacks from the ROAD dataset. Bechmarked methods include mean inter-message time, fitting a Gaussian curve, kernel density estimation, and binning. They found that the binning method outperformed the other methods, specially those fitting a distribution on inter-arrival times, in terms on AUC-ROC, AUC-PR, and F1 score. They reported a latency analysis for the binning algorithm.
@@ III-A Prior Work Closely Related to the Present Study
Pollicino et al. [40] introduced an experimental comparison on the performance evaluation of eight CAN IDS algorithms over two different datasets, an in-house one and one that is already publicly available (i.e., OTIDS [37]). The algorithms they bechmarked consume only time-based information from CAN frames as they are intended for detection of fabrication attacks. They open-sourced code implementations of these algorithms to allow reproducibility of research results and detail an unbiased experimental comparison.
@@
Compared to the benchmark studies mentioned above, the present paper is unique because it focuses on online IDS methods for masquerade attacks. Note that previous benchmark studies focused on: (1) comparing CAN IDS methods in an offline fashion (see works [21, 36, 35, 40, 18]); and (2) performing evaluation tests only on fabrication attacks (see works [21, 36, 38, 35, 39, 40]). By benchmarking CAN IDS methods on these two key aspects, we contribute to fill this gap of knowledge by analyzing the performance of unsupervised online CAN IDS in online fashion and testing them on the most stealthy CAN bus attacks, i.e., masquerade attacks. Our contribution enriches the state of the art by introducing a more realistic assessment of CAN IDS behavior in a realistic but constrained environment.
III-B Other Prior Work Related to the Present Study
Desta el al. [42] proposed a CAN IDS based on a LSTM sequence predictor model focused on predicting ID sequences. They performed in-house collection of CAN data using a real car. They replayed a CAN data log using the socketCAN API to evaluate their CAN IDS. The found that they can obtain better results by using the log loss of the predicted ID and the true ID.
@@ III-B Other Prior Work Related to the Present Study
Jedh et al. [44] evaluated four architecture designs for real-time CAN IDS under malicious speed reading message injections. They focus on a single ML-based IDS that captures the pattern of the sequences of CAN messages represented with a direct graph. They showed that the optimal architecture for deployment a real-time CAN IDS is based on two processes, a process for CAN monitoring and another one exclusively for anomaly detection.
@@
IV Methods
Common to all the methods benchmarked in this work is our focus on processing a set of time series representing signals in the CAN bus. For that, we assume that we have access to the time series representation obtained from CAN logs captured during a vehicle’s drive. Recall that the subsequent methods are expected to work in online fashion. This means that they process a time series stream to provide near real-time alerts.
Although the terms streaming and online algorithms are sometimes used interchangeably in the literature, we draw a distinction in line with prior surveys on data stream mining and adaptive learning [26, 47, 48]. In our context, streaming algorithms refer to those that process each data instance individually as it arrives, typically under strict memory and time constraints. In contrast, we use online algorithms to describe methods that process small batches or windows of data incrementally, allowing for limited buffering and adaptation over time. Throughout this work, we adopt the latter interpretation. When we refer to online algorithms, we mean algorithms designed to operate on continuously arriving data by processing small, sequential segments of the stream, updating models and decisions incrementally. Note that our inclination to benchmark unsupervised learning methods is guided in part by the fact that it is challenging to apply supervised methods on streaming data due to their static nature [49].
IV-A Time Windows
The presented benchmark is based on sliding time windows of the stream of time series data. The assumption behind using windows is that more recent information is more relevant for decision-making than past data. In particular, we focus on using sliding windows of fixed size for anomaly detection [50]. The windows are defined in terms of the number of observations or sequence-based windows [25]. This involves a sequence of $\omega$-length windows (i.e., $\omega\in\mathbb{Z}$) $W=\{W_{j}:j\geq 1\}$ with a sliding step (or offset $\delta\in[1,\omega]$) and observations arriving sequentially. Let $Y_{t}$ for $t\geq 0$ be a set of time series in a given window $W$, then $Y_{t}$ contains time series over the interval $[\delta\times(j-1),\delta\times(j-1)+\omega]$. Hence, $Y_{t}$ is a set of time series of the most recent $\omega$ observations up to time $t$ with $\delta$ observations that expire. Figure 2 shows this concept visually.
@@ IV-A Time Windows
IV-B Experimental Design
Here we describe our proposed benchmarking framework. More details can be found in Fig. 3. Our proposed framework has two phases: i) $P_{1}$: training phase and ii) $P_{2}$: testing phase. These phases have some sequential tasks. Some of them are unique to one phase while others are common to both phases. The tasks are: i) $T_{1}$: data collection and preprocessing, $T_{2}$: model fitting, and $T_{3}$: model testing. $P_{1}$ consists of tasks $T_{1}$ and $T_{2}$, whereas $P_{2}$ consists of tasks $T_{1}$, $T_{2}$, and $T_{3}$.
Figure 3: An overview of our proposed benchmark framework.
The proposed benchmark framework uses time windows for data processing and attack detection within the same time window. This means that our benchmark is focused on detecting attacks at the window level rather than at the message level. Note that both phases are considered to be accomplished in a light-weight device (e.g., Raspberry Pi or Nvidia Jetson Nano). We now explain the technical aspects of the framework in more detail.
IV-B1 Training Phase (P1)
@@
The first task common to all the algorithms in our benchmark is the configuration of the data collection and preprocessing pipeline. We assume that the benchmarked algorithms run on light-weight devices connected to the CAN bus through the OBD-II port or a dedicated ECU with direct access to the CAN bus. This setup allows continuous collection of raw CAN messages along with its translation from binary to time series using its corresponding decoding instructions. Note that there are available open-source CAN data loggers such as SavvyCAN and SocketCAN and commercial tools such as CANalyzer and VehicleSpy. We also assume that for translating CAN binary payloads to multidimensional time series, algorithms are loaded with OEM’s DBC file or use CAN decoding software such as LibreCAN, CAN-D, or CANMatch [51, 52, 53]. For a more comprehensive discussion on CAN signals reverse engineering, we refer the reader to the survey by Buscemi et al. [3].
@@
2) Model Fitting (T2):
We fit data-driven models for each of the benchmarked methods using the training data, i.e., $\mathbf{X}_{train}$. Fitted models in this task are then used as the regular (normal) state representation from which we derive comparisons with models fitted at the testing stage. The details of the fitted models for each of the becnhmarked methods are in Section IV-C.
IV-B2 Testing Phase (P2)
The testing phase executes similar tasks as already discussed in the training phase (see tasks T1 and T2 in Section IV-B1) and adds T3. Note that while P1 is run only once, the tasks in P2 are run continuously and sequentially as new messages are processed and there is a need to check for anomalies. We detail here changes with respect to specific parameters.

### v3

@@ Abstract
Vehicular controller area networks (CANs) are susceptible to masquerade attacks by malicious adversaries. In masquerade attacks, adversaries silence a targeted ID and then send malicious frames with forged content at the expected timing of benign frames. As masquerade attacks could seriously harm vehicle functionality and are the stealthiest attacks to detect in CAN, recent work has devoted attention to compare frameworks for detecting masquerade attacks in CAN. However, most existing works report offline evaluations using CAN logs already collected using simulations that do not comply with the domain’s real-time constraints. Here we contribute to advance the state of the art by presenting a comparative evaluation of four different non-deep learning (DL)-based unsupervised online intrusion detection systems (IDS) for masquerade attacks in CAN. Our approach differs from existing comparative evaluations in that we analyze the effect of controlling streaming data conditions in a sliding window setting. In doing so, we use realistic masquerade attacks being replayed from the ROAD dataset. We show that although evaluated IDS are not effective at detecting every attack type, the method that relies on detecting changes in the hierarchical structure of clusters of time series produces the best results at the expense of higher computational overhead. We discuss limitations, open challenges, and how the evaluated methods can be used for practical unsupervised online CAN IDS for masquerade attacks.
keywords:
Anomaly detection , controller area network , in-vehicle security , machine learning , online algorithms , streaming data.
1 Introduction
@@
IDS for detecting masquerade attacks in CAN are often based on unsupervised anomaly detection techniques aiming to characterize the regular relationships between physical signals in a vehicle [15, 18]. Some of them focus on mining time series relationships based on deep learning (DL) models that consequently have high computational cost [19, 16, 20]. In addition, most existing CAN IDS approaches are tested offline using datasets or CAN logs already collected from a real or simulated environment. This has been synthesized in existing comparative evaluation frameworks and studies in the CAN IDS space [21, 18]. The offline evaluation approach is different from online evaluation, in that the latter is performed using streaming CAN data from a vehicle, simulation, or data log replay [22]. Therefore, offline evaluation hinders the understanding of realistic capabilities and limitations of CAN IDS in a streaming environment. By conducting a comparative evaluation of non DL-based unsupervised online IDS for masquerade attacks in CAN, we seek to provide insights into the effectiveness and limitations of these methods in resource-constrained and streaming environments.
@@
In this paper, we address this gap of knowledge by introducing an empirical and unbiased comparative evaluation of four different non-DL-based unsupervised online IDS for masquerade attacks: Matrix Collelation Distribution, Matrix Correlation Correlation, Ganesan17 [23], and Moriano22 [15]. We focus on non DL-based approaches because vehicular data change over time and space. This dynamic nature makes it difficult for DL-based methods to extract relevant features, often leading to poor performance [24]. Here, Matrix Correlation Distribution and Matrix Correlation Correlation methods are simple proposed baselines while Ganesan17 and Moriano22 are already proposed methods in the literature. We test them in simulated masquerade attacks from the ROAD dataset [10]. We rely on a sliding window approach to simulate data log replays of a continuous data stream [25, 26]. Prior work has investigated how windowing strategies affect anomaly detection outcomes [27, 28], but has not explored overlapping configurations in multivariate streaming scenarios.
We analyze the joint effect of window size and offset on the AUC-ROC as evaluation metric for intrusion detection across the different attacks in the ROAD dataset. We also report an analysis of the testing time per window (TTW) [20] of CAN IDS as a proxy of inference time and discuss insights into the real-time detection capability of these approaches. We further contribute to advance the state of the art in this topic by making publicly available the code of the IDS considered in this study [29]. This furthers research in this topic by allowing other researchers to replicate our results and compare new algorithms with the evaluated IDS.
@@
Our results indicate that overall the evaluated unsupervised online CAN IDS are not effective in consistently detecting every attack type in the ROAD dataset. Attacks in the ROAD dataset have different levels of impact on the regular correlation patterns of signals. However, evaluation performance metrics tend to favor Moriano22 and Matrix Correlation Distribution methods. This aligns with the expectation that impacting correlations across signals tends to disrupt the hierarchy of their relationships and their pairwise similarity distribution. Notably, among the variety of attacks in the ROAD dataset, max_engine_attack and correlated_attack are the ones that can be detected more easily across CAN IDS. Our results also shed light on the impact that sliding window parameters (i.e., window length and offset) have on the detection task.
2 Background
2.1 CAN Protocol
@@ 2.3 CAN Vulnerabilities
  • 3.
    Masquerade attacks: This advanced attack requires two compromised ECUs, one as a strong attacker and one as a weak attacker. The weak attacker monitors the frequency of a targeted ID and, depending on the threat model, either compromises the software of the transmitting ECU or triggers a denial-of-service condition such as a bus-off attack to suppress its transmissions. Once the legitimate messages are suppressed, a strong attacker injects messages with the same ID at the original frequency. The ID frequency remains unchanged, but the source changes, and the messages contain forged content.
@@
3 Related Work
Here we review and contrast previous offline comparative evaluation studies on CAN IDS (Section 3.1) and online CAN IDS methods (Section 3.2).
3.1 Prior Work Closely Related to the Present Study
Ji et al. [21] presented a comparative analysis of four statistical-based CAN IDSes using entropy, clock skew, ID sequences, and CAN bus throughput. Compared methods relied only on ID timestamp arrivals. Their evaluation was done in simulated attack datasets. They found that the clock skew method performed best across attacks while the entropy method falls short on detecting replay attacks.
Dupont et al. [38] introduced a unifying framework for CAN IDS evaluation of time-based methods. In doing so, they used data from two live vehicles and a CAN prototype. They used both simulated attacks and publicly available datasets with attacks[39]. They found that among evaluated methods, they perform well only on attacks that generate significant disruptions in CAN traffic. They suggested that methods that consume bit representations and semantics of CAN messages will perform better.
Stachowski et al. [40] introduced a framework for evaluating CAN IDS. Their approach focused on the online evaluation of three existing solutions from identified vendors, so that CAN IDS are tested in real vehicles for real-time intrusion detection. Targeted ID attacks were introduced in test vehicles while vehicles were in motion and stationary. None of the tested CAN IDS was found effective for detecting the variety of attacks.
Blevins et al. [37] focused on evaluating time-based CAN IDS on the fuzzy and targetted ID attacks from the ROAD dataset. Compared methods include mean inter-message time, fitting a Gaussian curve, kernel density estimation, and binning. They found that the binning method outperformed the other methods, specially those fitting a distribution on inter-arrival times, in terms on AUC-ROC, AUC-PR, and F1 score. They reported a latency analysis for the binning algorithm.
@@ 3.1 Prior Work Closely Related to the Present Study
Pollicino et al. [42] introduced an experimental comparison on the performance evaluation of eight CAN IDS algorithms over two different datasets, an in-house one and one that is already publicly available (i.e., OTIDS [39]). The algorithms they evaluated consume only time-based information from CAN frames as they are intended for detection of fabrication attacks. They open-sourced code implementations of these algorithms to allow reproducibility of research results and detail an unbiased experimental comparison.
@@
Compared to the comparative evaluation studies mentioned above, the present paper is unique because it focuses on online IDS methods for masquerade attacks. Note that previous comparative evaluation studies focused on: (1) comparing CAN IDS methods in an offline fashion (see works [21, 38, 37, 42, 18]); and (2) performing evaluation tests only on fabrication attacks (see works [21, 38, 40, 37, 41, 42]). By evaluating CAN IDS methods on these two key aspects, we contribute to fill this gap of knowledge by analyzing the performance of unsupervised online CAN IDS in online fashion and testing them on the most stealthy CAN bus attacks, i.e., masquerade attacks. Our contribution enriches the state of the art by introducing a more realistic assessment of CAN IDS behavior in a realistic but constrained environment.
3.2 Other Prior Work Related to the Present Study
Desta el al. [44] proposed a CAN IDS based on a LSTM sequence predictor model focused on predicting ID sequences. They performed in-house collection of CAN data using a real car. They replayed a CAN data log using the socketCAN API to evaluate their CAN IDS. They found that they can obtain better results by using the log loss of the predicted ID and the true ID.
@@ 3.2 Other Prior Work Related to the Present Study
Jedh et al. [46] evaluated four architecture designs for real-time CAN IDS under malicious speed reading message injections. They focus on a single ML-based IDS that captures the pattern of the sequences of CAN messages represented with a direct graph. They showed that the optimal architecture for deploying a real-time CAN IDS is based on two processes, a process for CAN monitoring and another one exclusively for anomaly detection.
@@
Recent advances in unsupervised anomaly detection for tabular [49, 50] and image data [51] offer promising directions, but are not directly applicable to the multivariate streaming time series setting considered here.
4 Methods
Common to all the methods evaluated in this work is our focus on processing a set of time series representing signals in the CAN bus. For that, we assume that we have access to the time series representation obtained from CAN logs captured during a vehicle’s drive. Recall that the subsequent methods are expected to work in online fashion. This means that they process a time series stream to provide near real-time alerts.
Although the terms streaming and online algorithms are sometimes used interchangeably in the literature, we draw a distinction in line with prior surveys on data stream mining and adaptive learning [26, 52, 53]. In our context, streaming algorithms refer to those that process each data instance individually as it arrives, typically under strict memory and time constraints. In contrast, we use online algorithms to describe methods that process small batches or windows of data incrementally, allowing for limited buffering and adaptation over time. Throughout this work, we adopt the latter interpretation. When we refer to online algorithms, we mean algorithms designed to operate on continuously arriving data by processing small, sequential segments of the stream, updating models and decisions incrementally. Note that our inclination to comparative evaluation unsupervised learning methods is guided in part by the fact that it is challenging to apply supervised methods on streaming data due to their static nature [54].
Therefore, in this work, we follow Gama et al. [26, 52] and Avogadro et al. [53] in defining online anomaly detection as methods that operate on sequentially arriving data, processing it in small windows without requiring access to future observations. Our framework adheres to this setup and does not assume knowledge of future windows at inference time.
4.1 Time Windows
The presented comparative evaluation is based on sliding time windows of the stream of time series data. The assumption behind using windows is that more recent information is more relevant for decision-making than past data. In particular, we focus on using sliding windows of fixed size for anomaly detection [55]. The windows are defined in terms of the number of observations or sequence-based windows [25]. This involves a sequence of $\omega$-length windows (i.e., $\omega\in\mathbb{Z}$) $W=\{W_{j}:j\geq 1\}$ with a sliding step (or offset $\delta\in[1,\omega]$) and observations arriving sequentially. Let $Y_{t}$ for $t\geq 0$ be a set of time series in a given window $W$, then $Y_{t}$ contains time series over the interval $\delta\times(j-1),\delta\times(j-1)+\omega]$. Hence, $Y_{t}$ is a set of time series of the most recent $\omega$ observations up to time $t$ with $\delta$ observations that expire. Figure 2 shows this concept visually.
@@ 4.1 Time Windows
4.2 Experimental Design
Here we describe our proposed comparative evaluation framework. More details can be found in Fig. 3. Our proposed framework has two phases: i) $P_{1}$: training phase and ii) $P_{2}$: testing phase. These phases have some sequential tasks. Some of them are unique to one phase while others are common to both phases. The tasks are: i) $T_{1}$: data collection and preprocessing, $T_{2}$: representation extraction, and $T_{3}$: model testing. $P_{1}$ consists of tasks $T_{1}$ and $T_{2}$, whereas $P_{2}$ consists of tasks $T_{1}$, $T_{2}$, and $T_{3}$.
Figure 3: An overview of our proposed comparative evaluation framework.
The proposed comparative evaluation framework uses time windows for data processing and attack detection within the same time window. This means that our comparative evaluation is focused on detecting attacks at the window level rather than at the message level. Note that both phases can executed on edge-capable devices such as the Raspberry Pi or Nvidia Jetson Nano. While these are not automotive-grade embedded systems, they are widely used proxies in IDS research for evaluating feasibility under constrained compute and memory budgets [16]. The average processing times reported in Table 4 provide a proxy for lightweight deployability. We now explain the technical aspects of the framework in more detail.
4.2.1 Training Phase (P1)
@@
The first task common to all the algorithms in our comparative evaluation is the configuration of the data collection and preprocessing pipeline. We assume that the evaluated algorithms run on light-weight devices connected to the CAN bus through the OBD-II port or a dedicated ECU with direct access to the CAN bus. This setup allows continuous collection of raw CAN messages along with its translation from binary to time series using its corresponding decoding instructions. Note that there are available open-source CAN data loggers such as SavvyCAN and SocketCAN and commercial tools such as CANalyzer and VehicleSpy. We also assume that for translating CAN binary payloads to multidimensional time series, algorithms are loaded with OEM’s DBC file or use CAN decoding software such as LibreCAN, CAN-D, or CANMatch [56, 57, 58]. For a more comprehensive discussion on CAN signals reverse engineering, we refer the reader to the survey by Buscemi et al. [3].
@@
2) Representation Extraction (T2):
We fit data-driven models for each of the evaluated methods using the training data, i.e., $\mathbf{X}_{train}$. Fitted models in this task are then used as the regular (normal) state representation from which we derive comparisons with models fitted at the testing stage. The details of the fitted models for each of the becnhmarked methods are in Section 4.3.
4.2.2 Testing Phase (P2)
The testing phase executes similar tasks as already discussed in the training phase (see tasks T1 and T2 in Section 4.2.1) and adds T3. Note that while P1 is run only once, the tasks in P2 are run continuously and sequentially as new messages are processed and there is a need to check for anomalies. We detail here changes with respect to specific parameters.